4. A particle of mass in moves with momentumn of magnitude p.


(A) Show that the kinetic energy of the particle is KE = (P ^ 2)/(2m)

(B) Express the magnitude of the participle's momentum in terms of its kinetic energy and mass.


1
Expert's answer
2022-02-07T08:43:08-0500

The momentum of the particle by dfinition

"p=mv"

(A) The kinetic energy

"E_k=\frac{mv^2}{2}=\frac{m(p/m)^2}{2}=\frac{p^2}{2m}"

(B)

"p=\sqrt{2mE_k}"